# Recovering Parameters from Edge Fluctuations: Beta-Ensembles and Critically-Spiked Models

### v1

@@ 2.3.2. New Results
Theorem 1.2 refines current results on $\Lambda^{\theta}$’s rigidity in two directions: On the one hand, while the results of [29, 30] proved the existence of a single function $\mathcal{M}$ such that
(2.16) $\Lambda^{\theta}\cap B=\mathcal{M}(\Lambda^{\theta}\cap B^{c})\qquad\text{almost surely}$
@@
To the best of our knowledge, the optimality of number rigidity for $\Lambda^{\theta}$ (i.e., tolerance) has not been proved. That said, we expect that such a result could be proved using the technology developed in [26]. (The mutual absolute continuity of Palm measures of the Airy process—i.e., $\Lambda^{(1,2,\infty)}$—with the same number of fixed components proved in [13] serves as further evidence of this point.)
@@
If $\theta=(r,\beta,w)$ and $\tilde{\theta}=(\tilde{r},\tilde{\beta},\tilde{w})$ are such that $r_{0}+1/\beta\neq\tilde{r}_{0}+1/\tilde{\beta}$, then $\Lambda^{\theta}$ and $\Lambda^{\tilde{\theta}}$ are mutually singular.
@@ 2.3.3. Open Problems
Find a function $\Psi$ on the parameter space $\Theta$ such that $\Lambda^{\theta}$ and $\Lambda^{\tilde{\theta}}$ are mutually singular if and only if $\Psi(\theta)\neq\Psi(\tilde{\theta})$. Once $\Psi$ is found, characterize the possible relationships between the laws of $\Lambda^{\theta}$ and $\Lambda^{\tilde{\theta}}$ when $\Psi(\theta)=\Psi(\tilde{\theta})$ (e.g., equality, mutual absolute continuity, etc.).
@@ 2.4.1. Background
In [8, Page 2734], Bloemendal and Virág made the curious observation that any given spiked Wishart or Gaussian invariant model converges to an infinite number of SAOs with different parameters: Looking back at Theorem 2.13, if we add a trivial component to the spike $\ell=(\ell_{r-1},\ldots,\ell)$, thus transforming it into ~$\tilde{\ell}=(0,\ell_{r-1},\ldots,\ell_{0})$, then the matrix model remains unchanged. However, the parameter $r$ in the limit (2.14) increases by one, and the added component has a Dirichlet boundary condition (since $\ell_{r}=0$ corresponds to $w_{r+1}=\infty$ in (2.13)). If we combine this observation with the fact that the SAOs are obviously invariant with respect to permutations of their components (as their noises are induced by GOE, GUE, and GSE matrices), then we obtain the following from Theorem 2.13:
Corollary 2.25 ([8]).
@@ 2.4.2. New Result
More specifically, let $\tilde{\theta}=(\tilde{r},\beta,\tilde{w})$ be related to $\theta$ in such a way that (2.18) and (2.19) hold. Then, consider $\eta=(\kappa,\sigma,\upsilon)$ such that, similarly to the SAOs,
@@ 3. Trace Asymptotics
The covariance bound in Theorem 3.1 was proved in [29, (8.4)]. Consequently, we only need to prove (3.1). The remainder of this section is organized as follows: Before we prove (3.1), in Section 3.1 we show how Theorem 3.1 is used to construct the function $\mathcal{T}$ in (1.5), and thus prove Theorem 1.2. Then, in Section 3.2, we state a Feynman-Kac formula in proved in [29] for the expected trace $\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\right]$. Finally, in Sections 3.3 and 3.4, we use the Feynman-Kac formula in question to prove the asymptotics in (3.1).
@@
Let $c_{1},c_{2}>0$ be any fixed constants. Define the vanishing sequence of positive numbers $01\geq t_{1}>t_{2}>\cdots>$ 0 as
@@
$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\left(\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big{]}-\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big{]}\right]\right)=0\qquad\text{almost surely.}$.
@@
$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\mathrm{e}^{-t_{n}\Lambda^{\theta}_{k}/2}=1\qquad\text{almost surely.}$.
@@
The first step in our proof of (3.1) is to state the Feynman-Kac formula for $\mathbf{E}\left[\mathrm{Tr}\big{[}\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big{]}\right]$ obtained in [29, Theorem 4.9], which is what we will use to calculate the leading orders of the expected trace. This is stated as Lemma 3.13 below. Due to the complexity of the formula in question, we must set up a sizeable amount of notation to state it. First, we introduce the main stochastic process that drives the Feynman-Kac formula, which is the reflected Brownian bridge and its local times:
@@
We denote $B^{x}=(B|B(0)=x)$ and $B^{x,y}_{t}=(B|B(0)=x\text{ and }B(t)=y)$. Then, let $X=|B|$ be a reflected standard Brownian motion on $[0,\infty)$, and denote $X$’s transition kernel by
@@
If we were only dealing with scalar-valued Schrödinger operators on $[0,\infty)$, then these would be the only processes that are needed to state the Feynman-Kac formula. However, since we are dealing with the vector-valued setting, we need more tools. First, we set up some notation to describe the function spaces related to $\widehat{H}_{\theta,\eta}$:
@@
We denote $\mathcal{A}=\{1,\ldots,r\}\times[0,\infty)$, and we let $\mu$ denote the measure on $\mathcal{A}$ obtained by taking the product of the counting measure on $\{1,\ldots,r\}$ and the Lebesgue measure on $[0,\infty)$. Given $f:\mathcal{A}\to\mathbb{F}_{\beta}$, we denote
@@
Let $t>0$ and $x\geq 0$, let $n\geq 2$ be an even integer, and let $q\in\mathcal{P}_{n}$. Given a realization of $X^{x,x}_{t}$, we let $\mathrm{si}_{t,n,q,X^{x,x}_{t}}$ denote the unique random Borel probability measure on $[0,t)^{n}$ such that for every $s_{1},t_{1},\ldots,s_{n},t_{n}\in[0,t)$ with $s_{k}<t_{k}$ for $k=1,\ldots,n$, one has
@@ 6. $\bullet$
respectively denote $X^{x,x}_{t}$’s boundary local time at zero and local time collected on the intervals of $[0,t]$ where the jump process $\widehat{U}^{i}_{t}$ is equal to $j$. We note that since $[u,v)\mapsto\mathfrak{L}^{0}_{[u,v)}(X^{x,x}_{t})$ and $[u,v)\mapsto L^{y}_{[u,v)}(X^{x,x}_{t})$ induce measures on $[0,\infty)$, it is clear that
@@
We are now finally in a position to state our Feynman-Kac formula for the trace expectation, which is a special case of [29, Theorem 4.9]:
@@
Theorem 4.9 in [29] provides a formula for any joint moment of the form $\mathbf{E}\big{[}\mathrm{Tr}\big{[}\mathrm{e}^{-t_{1}\widehat{H}}\big{]}\cdots\mathrm{Tr}\big{[}\mathrm{e}^{-t_{n}\widehat{H}}\big{]}\big{]}$ for general vector-valued Schrödinger operators $\widehat{H}=-\frac{1}{2}\Delta+V+\xi$ on any one-dimensional domain $I\subset\mathbb{R}$. The statement in Lemma 3.13 can be recovered from the latter by taking $n=1$ and $t=t_{1}$, then considering what is called ”Case 2” in [29] (i.e., an operator on the half-line), and then taking the potential function $V(i,x)=\kappa x$.
@@ 3.3.3. Four or More Jumps
We now discuss $T_{\geq 4}(t)$. For this purpose, we begin with the observation that the combinatorial constant introduced in Definition 3.12 has a very simple upper bound (see [29, (4.2)]):
@@ 3.4.1. Preliminary: Brownian Motion Coupling
If we replace $F(X^{x,x}_{t})$ by $F(X^{x,x}_{t})\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}$, then (3.58) can be further simplified. Indeed, given that $B^{x,-x}_{t}$ must obviously reach zero as it travels from $x$ to $-x$, one has $\mathbf{1}_{\{\mathfrak{L}^{0}(|B^{x,-x}_{t}|)=0\}}=0$. Thus, in that case (3.58) simplifies to
@@
    By the strong Markov property and the symmetry of Brownian motion about 00, we notice that
@@
    (i.e., we can couple the two processes above by reflecting the path of $B^{x,x}_{t}$ after its first passage to zero, thus making it end at $-x$ instead of $x$).
@@
    By using the joint density of the running minimum and the current value of a Brownian motion (e.g., [54, Chapter III, Exercise 3.14]), we obtain that
@@
Finally, while (3.58) and (3.61) are useful for exact calculations (and thus necessary to obtain the explicit leading order terms in our asymptotics), simpler expressions can be used when we are only interested in upper estimates. For this purpose, we invoke [28, (5.9)], which states that for any nonnegative functional $F$, one has
@@
$\lim_{t\to 0^{+}}\|L_{1}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)\|_{2}^{2}=\lim_{t\to 0^{+}}\int_{0}^{\infty}L_{1}^{y-t^{-3/2}x}(B^{0,0}_{1})^{2}~{}\mathrm{d}y=\|L_{1}(B^{0,0}_{1})\|_{2}^{2}\qquad\text{almost surely.}$.
@@
By [52, (1)], we have the joint density
@@
then we get (3.75) (actually, the stronger statement RHS of (3.75) $=O(t^{3/2})$) by dominated convergence since $\mathrm{e}^{-\kappa t^{3/2}x}\uparrow 1$ as $t\to 0^{+}$. Toward this end, since $B^{0,0}_{1}(s)$ is Gaussian with mean zero and variance $s(1-s)$, a direct calculation yields
@@
Thanks to [52, (1)], for every $x,y\in\mathbb{R}$ and $\ell>0$, we have that
@@
Since $|B^{0,0}_{1}|$ is a Bessel bridge of dimension 1, the first expectation in the second line of (3.85) is finite thanks to the Gaussian tail bound for Bessel bridge maxima in [40, Remark 3.1]. The third expectation in the second line of (3.85) is finite thanks to [28, Lemma 5.8]. Finally, if we use the bound $(z_{1}+z_{2})^{2}\leq 2(z_{1}^{2}+z_{2}^{2})$, then we see that
@@
where the last inequality is from [28, Lemma 5.11]. In summary, (3.85) is finite, which means that we can now reduce (3.84) to
@@
thanks to the fact that $|B^{0,0}_{1}|$ supremum has Gaussian tails ([40, Remark 3.1]). Thirdly, it follows from (3.86) that
@@
Toward this end, noting that $\kappa t\mathrm{e}^{-\kappa tx}\mathrm{d}x$ is a probability measure on $[0,\infty)$ and that the square root function is concave, it follows from Jensen’s inequality that
@@
In the special case where $V(x)=x^{2}/2$, a classical calculation using a Selberg integral [55] yields the explicit formula $Z_{\beta,n}=Z_{G\beta E,n}$, where we define
@@
$-\tfrac{1}{n}\tfrac{\partial}{\partial\beta}\log Z_{G\beta E,n}=\tfrac{3}{8}n-\tfrac{1}{2}\log n-\tfrac{1}{4}\left(1+\log(\beta^{2}/4)-2\tfrac{\Gamma^{\prime}(1+\beta/2)}{\Gamma(1+\beta/2)}\right)+o(1)$
and $\tfrac{1}{n^{2}}\tfrac{\partial^{2}}{\partial^{2}\beta}\log Z_{G\beta E,n}=o(1)$. Combining this with (A.1), we get that when $V(x)=x^{2}/2$, the limit (2.4) holds. Using the well-known identity between the polygamma functions and the Hurwitz zeta function (e.g., [1, 6.4.10]), we have that for $\beta>0$,

### v2

@@ 2.3.2. New Results
Theorem 1.2 refines current results on $\Lambda^{\theta}$’s rigidity in two directions: On the one hand, while the results of [RigidityMSAO, RigiditySAO] proved the existence of a single function $\mathcal{M}$ such that
(2.16) $|\Lambda^{\theta}\cap B|=\mathcal{M}(\Lambda^{\theta}\cap B^{c})\qquad\text{almost surely}$
@@
To the best of our knowledge, the optimality of number rigidity for $\Lambda^{\theta}$ (i.e., tolerance) has not been proved. That said, we expect that such a result could be proved using the technology developed in [GangopadhyayGhoshTan]. (The mutual absolute continuity of Palm measures of the Airy process—i.e., $\Lambda^{(1,2,\infty)}$—with the same number of fixed components proved in [Bufetov18] serves as further evidence of this point.)
@@
[…]
@@ 2.3.3. Open Problems
[…]
@@ 2.4.1. Background
In [BloemendalVirag2, Page 2734], Bloemendal and Virág made the curious observation that any given spiked Wishart or Gaussian invariant model converges to an infinite number of SAOs with different parameters: Looking back at Theorem 2.13, if we add a trivial component to the spike $\ell=(\ell_{r-1},\ldots,\ell_{0})$, thus transforming it into $\tilde{\ell}=(0,\ell_{r-1},\ldots,\ell_{0})$, then the matrix model remains unchanged. However, the parameter $r$ in the limit (2.14) increases by one, and the added component has a Dirichlet boundary condition (since $\ell_{r}=0$ corresponds to $w_{r+1}=\infty$ in (2.13)). If we combine this observation with the fact that the SAOs are obviously invariant with respect to permutations of their components (as their noises are induced by GOE, GUE, and GSE matrices), then we obtain the following from Theorem 2.13:
Corollary 2.25 ([BloemendalVirag2]).
@@ 2.4.2. New Result
[…]
@@ 3. Trace Asymptotics
The covariance bound in Theorem 3.1 was proved in [RigidityMSAO, (8.4)]. Consequently, we only need to prove (3.1). The remainder of this section is organized as follows: Before we prove (3.1), in Section 3.1 we show how Theorem 3.1 is used to construct the function $\mathcal{T}$ in (1.5), and thus prove Theorem 1.2. Then, in Section 3.2, we state a Feynman-Kac formula proved in [RigidityMSAO] for the expected trace $\mathbf{E}\left[\mathrm{Tr}\big[\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big]\right]$. Finally, in Sections 3.3 and 3.4, we use the Feynman-Kac formula in question to prove the asymptotics in (3.1).
@@
Let $c_{1},c_{2}>0$ be any fixed constants. Define the vanishing sequence of positive numbers $1\geq t_{1}>t_{2}>\cdots>0$ as
@@
$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\left(\mathrm{Tr}\big[\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big]-\mathbf{E}\left[\mathrm{Tr}\big[\mathrm{e}^{-t_{n}\mathcal{H}_{\theta}/2}\big]\right]\right)=0\qquad\text{almost surely.}$
@@
$\lim_{m\to\infty}\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\mathrm{e}^{-t_{n}\Lambda^{\theta}_{k}/2}=1\qquad\text{almost surely.}$
@@
The first step in our proof of (3.1) is to state the Feynman-Kac formula for $\mathbf{E}\left[\mathrm{Tr}\big[\mathrm{e}^{-t\widehat{H}_{\theta,\eta}}\big]\right]$ obtained in [RigidityMSAO, Theorem 4.9], which is what we will use to calculate the leading orders of the expected trace. This is stated as Lemma 3.13 below. Due to the complexity of the formula in question, we must set up a sizeable amount of notation to state it. First, we introduce the main stochastic process that drives the Feynman-Kac formula, which is the reflected Brownian bridge and its local times:
@@
We denote $B^{x}=(B|B(0)=x)$ and $B^{x,y}_{t}=(B|B(0)=x\text{ and }B(t)=y)$. Then, let $X=|B|$ be a reflected standard Brownian motion on $0,\infty)$, and denote $X$’s transition kernel by
@@
If we were only dealing with scalar-valued Schrödinger operators on $0,\infty)$, then these would be the only processes that are needed to state the Feynman-Kac formula. However, since we are dealing with the vector-valued setting, we need more tools. First, we set up some notation to describe the function spaces related to $\widehat{H}_{\theta,\eta}$:
@@
We denote $\mathcal{A}=\{1,\ldots,r\}\times[0,\infty)$, and we let $\mu$ denote the measure on $\mathcal{A}$ obtained by taking the product of the counting measure on $\{1,\ldots,r\}$ and the Lebesgue measure on $0,\infty)$. Given $f:\mathcal{A}\to\mathbb{F}_{\beta}$, we denote
@@
Let $t>0$ and $x\geq 0$, let $n\geq 2$ be an even integer, and let $q\in\mathcal{P}_{n}$. Given a realization of $X^{x,x}_{t}$, we let $\mathrm{si}_{t,n,q,X^{x,x}_{t}}$ denote the unique random Borel probability measure on $0,t)^{n}$ such that for every $s_{1},t_{1},\ldots,s_{n},t_{n}\in[0,t)$ with $s_{k}<t_{k}$ for $k=1,\ldots,n$, one has
@@ 6. $\bullet$
respectively denote $X^{x,x}_{t}$’s boundary local time at zero and local time collected on the intervals of $[0,t]$ where the jump process $\widehat{U}^{i}_{t}$ is equal to $j$. We note that since $u,v)\mapsto\mathfrak{L}^{0}_{[u,v)}(X^{x,x}_{t})$ and $u,v)\mapsto L^{y}_{[u,v)}(X^{x,x}_{t})$ induce measures on $0,\infty)$, it is clear that
@@
We are now finally in a position to state our Feynman-Kac formula for the trace expectation, which is a special case of [RigidityMSAO, Theorem 4.9]:
@@
Theorem 4.9 in [RigidityMSAO] provides a formula for any joint moment of the form $\mathbf{E}\big[\mathrm{Tr}\big[\mathrm{e}^{-t_{1}\widehat{H}}\big]\cdots\mathrm{Tr}\big[\mathrm{e}^{-t_{n}\widehat{H}}\big$$\big]$ for general vector-valued Schrödinger operators $\widehat{H}=-\frac{1}{2}\Delta+V+\xi$ on any one-dimensional domain $I\subset\mathbb{R}$. The statement in Lemma 3.13 can be recovered from the latter by taking $n=1$ and $t=t_{1}$, then considering what is called ”Case 2” in [RigidityMSAO] (i.e., an operator on the half-line), and then taking the potential function $V(i,x)=\kappa x$.
@@ 3.3.3. Four or More Jumps
We now discuss $T_{\geq 4}(t)$. For this purpose, we begin with the observation that the combinatorial constant introduced in Definition 3.12 has a very simple upper bound (see [RigidityMSAO, (4.2)]):
@@ 3.4.1. Preliminary: Brownian Motion Coupling
If we replace $F(X^{x,x}_{t})$ by $F(X^{x,x}_{t})\mathbf{1}_{\{\mathfrak{L}^{0}_{t}(X^{x,x}_{t})=0\}}$, then (3.58) can be further simplified. Indeed, given that $B^{x,-x}_{t}$ must obviously reach zero as it travels from $x$ to x-x, one has $\mathbf{1}_{\{\mathfrak{L}^{0}(|B^{x,-x}_{t}|)=0\}}=0$. Thus, in that case (3.58) simplifies to
@@
    By the strong Markov property and the symmetry of Brownian motion about 0, we notice that
@@
    (i.e., we can couple the two processes above by reflecting the path of $B^{x,x}_{t}$ after its first passage to zero, thus making it end at x-x instead of $x$).
@@
    By using the joint density of the running minimum and the current value of a Brownian motion (e.g., [RevuzYor, Chapter III, Exercise 3.14]), we obtain that
@@
Finally, while (3.58) and (3.61) are useful for exact calculations (and thus necessary to obtain the explicit leading order terms in our asymptotics), simpler expressions can be used when we are only interested in upper estimates. For this purpose, we invoke [GaudreauLamarreEJP, (5.9)], which states that for any nonnegative functional $F$, one has
@@
$\lim_{t\to 0^{+}}\|L_{1}(|B^{t^{-3/2}x,t^{-3/2}x}_{1}|)\|_{2}^{2}=\lim_{t\to 0^{+}}\int_{0}^{\infty}L_{1}^{y-t^{-3/2}x}(B^{0,0}_{1})^{2}~\mathrm{d}y=\|L_{1}(B^{0,0}_{1})\|_{2}^{2}\qquad\text{almost surely.}$
@@
By [Pitman, (1)], we have the joint density
@@
then we get (3.75) (actually, the stronger statement RHS of (3.75) $O(t^{3/2})$) by dominated convergence since $\mathrm{e}^{-\kappa t^{3/2}x}\uparrow 1$ as $t\to 0^{+}$. Toward this end, since $B^{0,0}_{1}(s)$ is Gaussian with mean zero and variance s(1s)s(1-s), a direct calculation yields
@@
Thanks to [Pitman, (1)], for every $x,y\in\mathbb{R}$ and $\ell>0$, we have that
@@
Since $|B^{0,0}_{1}|$ is a Bessel bridge of dimension 1, the first expectation in the second line of (3.85) is finite thanks to the Gaussian tail bound for Bessel bridge maxima in [GruetShi, Remark 3.1]. The third expectation in the second line of (3.85) is finite thanks to [GaudreauLamarreEJP, Lemma 5.8]. Finally, if we use the bound $(z_{1}+z_{2})^{2}\leq 2(z_{1}^{2}+z_{2}^{2})$, then we see that
@@
where the last inequality is from [GaudreauLamarreEJP, Lemma 5.11]. In summary, (3.85) is finite, which means that we can now reduce (3.84) to
@@
thanks to the fact that $|B^{0,0}_{1}|$ supremum has Gaussian tails ([GruetShi, Remark 3.1]). Thirdly, it follows from (3.86) that
@@
Toward this end, noting that $\kappa t\mathrm{e}^{-\kappa tx}\mathrm{d}x$ is a probability measure on $0,\infty)$ and that the square root function is concave, it follows from Jensen’s inequality that
@@
In the special case where $V(x)=x^{2}/2$, a classical calculation using a Selberg integral [Selberg] yields the explicit formula $Z_{\beta,n}=Z_{G\beta E,n}$, where we define
@@
$\tfrac{1}{n}\tfrac{\partial}{\partial\beta}\log Z_{G\beta E,n}=\tfrac{3}{8}n-\tfrac{1}{2}\log n-\tfrac{1}{4}\left(1+\log(\beta^{2}/4)-2\tfrac{\Gamma^{\prime}(1+\beta/2)}{\Gamma(1+\beta/2)}\right)+o(1)$
and $\tfrac{1}{n^{2}}\tfrac{\partial^{2}}{\partial^{2}\beta}\log Z_{G\beta E,n}=o(1)$. Combining this with (A.1), we get that when $V(x)=x^{2}/2$, the limit (2.4) holds. Using the well-known identity between the polygamma functions and the Hurwitz zeta function (e.g., [AbramowitzStegun, 6.4.10]), we have that for $\beta>0$,